OpenStudy (anonymous):

I need help solving this system of linear equations. 6x+2y=6 18x+13y=60

OpenStudy (anonymous):

y=6 and x =-1

OpenStudy (anonymous):

Get em girlll heyyy:D

OpenStudy (amistre64):

divide a 3 out of the bottom

OpenStudy (anonymous):

Can you explain how to solve this? Pleaseeee :D

OpenStudy (anonymous):

now multiply the first equation by 3 which makes it 18x +6y =18 18x +13y=60 7y=42 y=6 and x = -6/6=-1

OpenStudy (amistre64):

6x+ 2y=6 18x+13y=60 ; /-3, -3 will be quicker maybe 6x+2(3)/3 y = 6 -6x - 13 /3 y = -20 ----------------- -7/3 y = -14; solve for y, * -3/7 y = -14(-3)/7 = 2(3) = 6

OpenStudy (amistre64):

i was hoping the bottom equation would be easier lol ... same results just had to deal with the fraction